OpenStudy (anonymous):

A square and a rectangle have the same perimeter. The length of one side of the square is 11cm and the longer side of the rectangle is 13cm. Find the shorter side of the rectangle.

OpenStudy (ajprincess):

Perimeter=2(length+width) Since all the sides of a square are equal Perimeter=2(11+11) =44 For rectangle 44=2(13+w) 44=26+2w 18=2w w=9

OpenStudy (ajprincess):

Is that clear?

OpenStudy (anonymous):

no i dont get it.

OpenStudy (ajprincess):

The length of one side of a square is given as11. so the perimeter will be 11+11+11+11 which is 44. Is t clear upto this point?

OpenStudy (ajprincess):

since all the sides of the square are equal.